OpenStudy (anonymous):

(4 over 3) x(2 over 3) + = 6

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[\frac{ 4 }{ 3 }x+\frac{ 2 }{ 3 }=6\]

OpenStudy (anonymous):

multiply everything by 3, that should help you a lot to start

OpenStudy (nurali):

4/3x+2/3=6 4/3x=6-2/3 4/3x=16/3 x=16*3/3*4 x=4